ROTATIONAL DEPENDENCE OF INTRAMOLECULAR DYNAMICS IN ACETYLENE AT LOW VIBRA-TIONAL EXCITATION AS DEDUCED FROM HIGH RESOLUTION SPECTROSCOPY

\underline{DAVID S. PERRY}, ANTHONY MILLER, Department of Chemistry, The University of Akron, Akron OH 44325-3601; B. AMYAY, A. FAYT, and M. HERMAN, Laboratoire de Chimie Quantique et Photophysique, Université libre de Bruxelles, B-1050 Brussels, Belgium.

The link between energy-resolved spectra and time-resolved dynamics is explored quantitatively for acetylene (^{12}C_{2}H_{2}), X̃^{1}Σ_{g}^{+} with up to 8,600 cm^{−1}of vibrational energy. This comparison is based on the extensive knowledge of the vibrationrotation energy levels and on the model Hamiltonian used to fit them to high precision.^{a} Simulated intensity borrowing features in high resolution absorption spectra and predicted survival probabilities for intramolecular vibrational redistribution (IVR) are first investigated for the ν_{4} + ν_{5} and ν_{3} bright states, for J = 2, 30 and 100. The dependence of the results on the rotational quantum number and on the choice of vibrational bright state reflects the interplay of three kinds of off-diagonal resonances: anharmonic, rotational l−type, and Coriolis. The dynamical quantities used to characterize the calculated time-dependent dynamics are the dilution factor ϕ_{d}, the IVR lifetime τ_{IV R}, and the recurrence time τ_{rec}. For the two bright states ν_{3} +2ν_{4} and 7ν_{4}, the collisionless dynamics for thermally averaged rotational distributions at T = 27, 270 and 500 K were calculated from the available spectroscopic data. For the 7ν_{4} bright state, an apparent irreversible decay of is found. In all cases, the model Hamiltonian allows a detailed calculation of the energy flow among all of the coupled zeroth-order vibration-rotation states.

aB. Amyay, S. Robert, M. Herman, A. Fayt, B. Raghavendra, A. Moudens, J. Thivin, B. Rowe, and R. Georges, J. Chem. Phys.

\underline{131}, 114301 (2009).